OpenStudy (anonymous):

Help please will fan and give a medal. 2y^4 over 9x^5y divided y over 6x^2

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

4y^2 over 3x^3

OpenStudy (ikram002p):

\(\frac{2y^4}{9x^5y}/ \frac{y}{6x^2} \) first thing to do is change it to product instead of divide like this \(\frac{2y^4}{9x^5y}. \frac{6x^2}{y}\)

OpenStudy (anonymous):

Just go here to help you with the steps http://www.cymath.com/answer.php?q=(2y%5E4)%2F(9x%5E5y)*(6x%5E2)%2F(y)

OpenStudy (ikram002p):

http://prntscr.com/520kix

OpenStudy (ikram002p):

http://prntscr.com/520ktc

OpenStudy (ikram002p):

http://prntscr.com/520l5w

OpenStudy (ikram002p):

http://prntscr.com/520lnr

OpenStudy (ikram002p):

:)

OpenStudy (anonymous):

So is correct answer 4/3y^2x^3

OpenStudy (anonymous):

yes that the answer

OpenStudy (ikram002p):

nope its 4y^2/3x^3

OpenStudy (anonymous):

4y^2 over 3x^3

OpenStudy (anonymous):

Ok thank you both @ikram002p and @Ese_Muerte